### v1

# The $B^{+}\to K^{+}\nu\bar{\nu}$ decay as a search for the QCD axion

Abstract
We reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of the decay $B^{+}\to K^{+}a$, where $a$ denotes an axion or axion-like particle. This is possible in the kinematic regions where the di-neutrino invariant mass squared, $q^{2}$, can be identified with the assumed $m_{a}^{2}$. While $q^{2}$ is reconstructed only in the Hadronic Tag Analysis (HTA), it is not directly accessible in the Inclusive Tag Analysis (ITA), which provides much larger statistics and hence higher sensitivity. In ITA, $q^{2}$ is proxied by a reconstructed variable, $q^{2}_{\text{rec}}$, usually modeled through Belle II simulations. We show that the mapping between $q^{2}$ and $q^{2}_{\text{rec}}$ can be accurately derived from kinematic arguments alone—without relying on internal experimental inputs. Using this relation and publicly available efficiencies, we obtain the strongest existing bounds on the coupling-rescaled Peccei-Quinn scale, $|(F_{V})_{sb}|\geq 0.9\times 10^{9}$ GeV, improving the latest bound in the literature by an order of magnitude. We further show that the bound depends only marginally on the assumed value of the di-neutrino signal strength $\mu_{\nu\bar{\nu}}$, whether fixed or floated. This establishes $B^{+}\to K^{+}\nu\bar{\nu}$ as a double probe—of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, and of new light, elusive particles produced via $B^{+}\to K^{+}a$—the two probes working independently to an excellent approximation.
preprint: LAPTH-036/25
Introduction— Spin-zero, odd-parity particles below the GeV scale—collectively known as axions or axion-like particles (ALPs)—are hypothetical states actively sought through laboratory experiments [1, 2, 3], astrophysical data [4, 5, 6], cosmological probes [7], and theoretical studies [8, 9, 10]. A central focus of this research domain is the “invisible” [11, 12, 13, 14] QCD axion [15, 16, 17, 18], henceforth the axion, which elegantly solves the strong-$CP$ problem—the unexplained suppression of a $CP$-violating QCD term. Given the extreme suppression, $O(10^{-10})$, neither coincidence nor anthropic reasoning provides a satisfactory explanation [19, 20]. The axion also offers a viable Dark Matter candidate [21, 22, 23]. Unlike generic ALPs, its mass $m_{a}$ and decay constant $f_{a}$ are linked: smaller $m_{a}$ implies weaker couplings, scaling as $1/f_{a}$.
Existing bounds indicate that $m_{a}$ must be several orders of magnitude below the pion mass [9]. Yet this small mass also opens the possibility that axions could be copiously produced at high-intensity colliders, without having to resort to high energies. One can consider decays of $K$, $D$, and $B$ mesons, or even the baryonic counterparts, decaying to fully reconstructible final states, plus the axion, which is assumed to escape without decay nor interaction within the detector volume. This represents the most general and conservative scenario.
Since the QCD axion manifests as an energy and momentum imbalance—equal in magnitude, as the axion has negligible mass with respect to the lightest physical scale of the decay—it is crucial that the remaining final-state particles be reconstructible with high precision. Moreover, the identity and momentum of the decaying hadron must be accurately known to better control the many backgrounds involving elusive particles that can mimic the axion.
Belle II [24] is particularly well positioned to satisfy both of these constraints. It collides $e^{-}$ and $e^{+}$ beams with momenta in approximately opposite directions and magnitudes of $7$ and $4$ GeV  [^1] The spread, or bite, around these momenta is tiny and completely negligible for the purposes of this paper., respectively. The center-of-mass energy corresponds to the $b\bar{b}$ resonance known as $\Upsilon(4S)$, which subsequently decays with near-unit probability into charged or neutral $B\bar{B}$ meson pairs. Signals are then studied in the decays of one of the produced mesons, $B_{\mathrm{s}ig}$, while the other meson is often used as a tag decay—allowing the kinematics of the signal decay to be fully reconstructed and the flavour of $B_{\mathrm{s}ig}$ to be identified.
In this paper we will consider the $B^{+}\to K^{+}a$ decay  [^2] The conjugated mode with negatively-charged mesons is assumed to be included throughout our study.. The axion-hadron interaction responsible for the decay inherits in a calculable way from the underlying axion-quark interaction, which just below the electroweak scale can be written most generally as [27]
$\mathcal{L}_{a}\supset\;$ $\frac{\partial_{\mu}a}{2f_{a}}\sum_{i,j=1}^{3}\bar{d}_{i}\gamma^{\mu}\bigl[(k_{V})_{d_{i}d_{j}}+\gamma_{5}(k_{A})_{d_{i}d_{j}}\bigl]d_{j}+\left\{d\to u\right\}$
$-\frac{\alpha_{s}}{8\pi}\frac{a}{f_{a}}G\tilde{G}~,$ (1)
where $u_{i},d_{i}$ denote up- and down-type quarks of generation $i$ in the mass eigenbasis. The $G\tilde{G}$ operator  [^3] The sign difference in the $f_{a}$-defining $aG\tilde{G}$ term in Eq.The $B^{+}\to K^{+}\nu\bar{\nu}$ decay as a search for the QCD axion with respect to Ref.[9] arises from our use of opposite $\epsilon$-tensor conventions. can be traded for an axion-field dependent quark mass matrix by redefining the quark fields through $q\to\exp[i(Q_{V}+Q_{A}\gamma_{5})\,a/(2f_{a})]q$, with $Tr Q_{A}=1$ [27]. Although the hermitian matrices $Q_{V},Q_{A}$ can be chosen as diagonal, they introduce an unphysical reparametrization ambiguity [27].
The process $B^{+}\to K^{+}a$, on which this paper is focused, involves the complex coupling $(k_{V})_{sb}$. The calculation of this process requires the following hadronic matrix element
$\langle K^{-}(k)|\bar{s}b|B^{-}(p)\rangle=\frac{m_{B}^{2}-m_{K}^{2}}{m_{b}-m_{s}}f_{0}(m_{a}^{2})~.$ (2)
Here, form-factor values and conventions are taken from Ref. [29]. The relation in Eq. (2) imply the following branching ratios (BRs) [30, 31]
$\mathcal{B}(B^{-}\to K^{-}a)=\frac{\tau_{B}|(k_{V})_{sb}|^{2}\,|\vec{k}|m_{B}^{2}}{32\pi f_{a}^{2}}\left(1-\frac{m_{K}^{2}}{m_{B}^{2}}\right)^{2}f_{0}^{2}(m_{a}^{2})~,$ (3)
where $|\vec{k}|=\lambda^{1/2}(m_{B}^{2},m_{K^{(*)}}^{2},m_{a}^{2})/(2m_{B})$ and $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$. While the theoretically well-motivated QCD axion requires $m_{a}\simeq 0$, we will also discuss the general case of an ALP. Here and henceforth, the axion-quark couplings are understood at a scale close to $m_{B}$.
We search for the signal Eq. (3) in data collected by Belle II for the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements. Belle II employs two nearly uncorrelated datasets, analyzed using the “classic” Hadronic Tagging Analysis (HTA) method and the more recent Inclusive Tagging Analysis (ITA) method, respectively [32, 33, 34]. The ITA method achieves much larger statistics and thus dominates the sensitivity to the measurement. Reliably recasting this dataset for our search thus hinges on one key piece of non-public information: the mapping between the true di-neutrino invariant mass squared, $q^{2}$, which is inaccessible within the ITA method, and its reconstructed counterpart, $q^{2}_{{\mathrm{r}ec}}$, defined below. We show that the $q^{2}$ distribution as a function of $q^{2}_{{\mathrm{r}ec}}$ can be reliably inferred from kinematic arguments alone—i.e., without requiring internal information from the experiment. Using this mapping and publicly available efficiencies, we demonstrate that this search yields the strongest existing bounds on the magnitudes of the fundamental QCD-axion coupling $(k_{V})_{sb}$.
Both of the aforementioned ITA and HTA methods yield a central value in excess of the SM prediction [35, 36]. This excess motivates the following parameterization of the BR for $B^{+}\to K^{+}\nu\bar{\nu}$—differential in the squared di-neutrino invariant mass $q^{2}$
$d\mathcal{B}_{\nu\bar{\nu}}(\mu)/dq^{2}~\equiv~\mu_{\nu\bar{\nu}}\,d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}/dq^{2}~,$ (4)
where on the r.h.s. we use the SM prediction [35, 37, 38], rescaled by a signal-strength factor. This rescaling allows one to disentangle the presence of new effects in the $B^{+}\to K^{+}\nu\bar{\nu}$ mode from those due to the two-body decay $B^{+}\to K^{+}a$, and to test whether the bound inferred on the $(k_{V})_{sb}$ coupling is affected by possible new-physics contamination in the $B^{+}\to K^{+}\nu\bar{\nu}$ sample.
Within the ITA, Belle II cannot directly reconstruct $q^{2}\equiv m_{\nu\bar{\nu}}^{2}=(p-k)^{2}$ and thus constructs a proxy variable, $q^{2}_{{\mathrm{r}ec}}$ [32], in terms of the known $s\equiv(p_{e^{-}}+p_{e^{+}})^{2}$ and the energy of the final-state kaon in the $B\bar{B}$ rest frame (hereafter denoted with a star). In this frame, the signal $B$-meson momentum is known in magnitude, $|\vec{p}_{B}^{\,*}|=0.3298~{\mathrm{G}eV}$, and is distributed isotropically. In turn, in the $B$ rest frame (denoted with a double star), the kaon is emitted isotropically. One can therefore write
$\begin{split}q^{2}_{{\mathrm{r}ec}}~\equiv~&\frac{s}{4}+m_{K}^{2}-\sqrt{s}E_{K}^{*}~=~q^{2}+E_{B}^{*2}-m_{B}^{2}-2\,\vec{p}_{B}^{\,*}\cdot\vec{p}_{K}^{\,*}(q^{2})\\ ~=~&q^{2}+E_{B}^{*2}-m_{B}^{2}-2|\vec{p}_{B}^{\,*}|\gamma^{**}\left(\beta^{**}E_{K}^{**}(q^{2})+|\vec{p}_{K}^{\,**}(q^{2})|\cos\theta^{**}\right)~,\end{split}$ (5)
where $|\vec{p}_{K}^{\,**}|=\lambda^{1/2}(m_{B}^{2},m_{K}^{2},q^{2})/(2m_{B})$ and $\beta^{**}=|\vec{p}_{B}^{\,*}|/E_{B}^{*}=0.06234$ is the boost parameter in the Lorentz transformation from the $B$ rest frame to the $B\bar{B}$ rest frame. Finally, $\theta^{**}$ is the polar angle of the kaon momentum in the $B$-meson rest frame relative to the $B$-meson momentum in the $B\bar{B}$ rest frame. Importantly, $\theta^{**}$ constitutes the only unmeasured quantity in Eq. (5). The distribution of $q^{2}_{{\mathrm{r}ec}}$ for a given $q^{2}$ can therefore be derived from Eq. (5) with $c^{**}\equiv\cos\theta^{**}\in[-1,1]$ uniformly distributed. The sub-percent-level uncertainties [39] smearing the energies and momenta that appear in Eq. (5) have a completely negligible impact on the resulting $q^{2}(q^{2}_{{\mathrm{r}ec}})$ distribution.
The physical branching ratios and the observed yields are then related in the ITA as
$\frac{dN_{X}}{dq^{2}_{{\mathrm{r}ec}}}~=~N_{B}\int dq^{2}\int dc^{**}\,\frac{1}{2}\,\delta\Bigl(q^{2}_{{\mathrm{r}ec}}-q^{2}_{{\mathrm{r}ec}}(q^{2},c^{**})\Bigl)\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~\equiv~N_{B}\int dq^{2}\,f_{q^{2}_{{\mathrm{r}ec}}}(q^{2})\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~,$ (6)
where $N_{B}=387(6)\times 10^{6}$ [32]  [^4] The below-2% $N_{B}$ error has a negligible impact on our analysis, because $n_{\nu\bar{\nu},{\mathrm{b}kg}}^{i}$ have an error approaching 10%. is the number of $B$ candidates in the given decay channel $X$, $\mathcal{B}$ the corresponding BR, and $\epsilon_{\text{ITA}}(q^{2})$ the ITA-specific selection efficiency as a function of $q^{2}$, which can be obtained from Ref. [32]. The function $q^{2}_{{\mathrm{r}ec}}(q^{2},c^{**})$ appearing in the delta function is the r.h.s. of Eq. (5). In the last equality we make contact with the smearing function $f_{q^{2}_{{\mathrm{r}ec}}}$ [41, 31, 42], which is typically assumed to require Belle II simulation.
The $q^{2}$ vs. $q^{2}_{{\mathrm{r}ec}}$ correlation implied by Eq. (6) is shown in Fig. 1.
Figure 1: Correlation between the true ($q^{2}$) and the reconstructed ($q^{2}_{{\mathrm{r}ec}}$) di-neutrino invariant mass squared. The plot is normalised to $10^{5}$ observed events.
This correlation plot may be compared with figure 5.1 of Ref. [43], which is the output of internal Belle II simulations. The two figures display good agreement, aside from an inconsequential low-$q^{2}$ tail visible in figure 5.1 of Ref. [43]  [^5] This tail originates from radiative decays with undetected photons. Strictly speaking, $B\to K\nu\bar{\nu}(+n\gamma)$ is a distinct process from $B\to K\nu\bar{\nu}$ and should be subtracted if sizeable, as in $B_{s}\to\mu^{+}\mu^{-}$[50]. Here, however, the effect is negligible (well below 1% in Ref. [43]), since the lightest emitter is the $K^{+}$, which makes the exponent $b$ in ${\mathcal{B}}(B\to K\nu\bar{\nu}(+n\gamma))=\Bigl(\tfrac{2\Delta E}{m_{B}}\Bigr)^{\tfrac{\alpha}{4\pi}b}{\mathcal{B}}(B\to K\nu\bar{\nu})$ parametrically small. Here $\Delta E$ denotes the minimal photon energy in the $B$ rest frame required for detectability—a “theoretical” scale usually proxied by the $\sqrt{q^{2}}$ resolution. .
To make the origin of the $q^{2}_{{\mathrm{r}ec}}$ smearing explicit, we recall that the relation between $q^{2}$ and $q^{2}_{{\mathrm{r}ec}}$, Eq. (5), introduces a uniformly distributed variable from an unmeasured angle. This kinematic unknown constitutes an irreducible source of uncertainty, which is not immediately transparent in simulation output. The good agreement between Fig. 1 and figure 5.1 of Ref. [43] shows that this mechanism alone accounts for the observed correlation, with any additional effects being negligible.
Within the HTA, instead, Belle II can directly reconstruct the $q^{2}$ of the events, so the physical BRs and observed yields are simply related as
$\frac{dN_{X}}{dq^{2}}~=~N_{B}\,\epsilon_{\text{HTA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~,$ (7)
where $\epsilon_{\text{HTA}}(q^{2})$ is the HTA-specific selection efficiency as a function of $q^{2}$, also available in Ref. [32].
Analysis— To probe the $B^{+}\to K^{+}a$ decay through the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements at Belle II, we construct a likelihood ratio test statistic based solely on the public data of Fig. 17 (ITA) and Fig. 20 (HTA) of Ref. [32]. The likelihood function $\mathcal{L}$ is constructed as the product of two contributions, i.e. ${\mathcal{L}}={\mathcal{L}}_{1}{\mathcal{L}}_{2}$.
The first and main term in the likelihood function represents a Poisson distribution $P$ for each bin, labeled by an index $i$, assuming the bins to be uncorrelated: ${\mathcal{L}}_{1}=\prod_{i}^{n_{\mathrm{b}in}}P(n^{i}_{\mathrm{o}bs};\lambda^{i})$. Here $n^{i}_{\mathrm{o}bs}$ is the number of observed events in bin $i$, defined by $q^{2}\in[q^{2}_{i},q^{2}_{i+1})$, and $\lambda^{i}$ is the corresponding expected number of events, which depends on $\mathcal{B}(B^{+}\to K^{+}a)\equiv\mathcal{B}_{a}$ and on the signal strength $\mu_{\nu\bar{\nu}}$—the two key quantities to be determined:
$\begin{split}\lambda^{i}(\mathcal{B}_{a},\mu_{\nu\bar{\nu}})~=~\mathcal{B}_{a}\,n^{i}_{a}+\mu_{\nu\bar{\nu}}\,n^{i}_{\nu\bar{\nu}}+(1+\sigma^{\text{bkg}}\,\theta)\,n^{i}_{\text{bkg}}~.\end{split}$ (8)
This expression applies to either the ITA or HTA dataset, where $n^{i}_{a}$, $n^{i}_{\nu\bar{\nu}}$, and $n^{i}_{\mathrm{b}kg}$ are the yields in the $i^{\mathrm{t}h}$ bin for the signal, the $B^{+}\to K^{+}\nu\bar{\nu}$ decay, and the background, respectively. Within the two methods, the first two components on the r.h.s. of Eq. (8) are defined as:
$n^{i}_{a,\text{ITA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}_{{\mathrm{r}ec}}\,f_{q^{2}_{{\mathrm{r}ec}}}(m_{a}^{2})\,\epsilon_{\text{ITA}}(m_{a}^{2})~,$ (9)
$n^{i}_{\nu\bar{\nu},\text{ITA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}_{{\mathrm{r}ec}}\int dq^{2}\,f_{q^{2}_{{\mathrm{r}ec}}}(q^{2})\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}}{dq^{2}}~,$ (10)
$n^{i}_{a,\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(q^{2})\,g_{a}(q^{2})~,$ (11)
$n^{i}_{\nu\bar{\nu},\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(q^{2})\,\frac{d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}}{dq^{2}}~.$ (12)
These expressions can be understood in terms of Eqs. (6)-(7), bearing in mind that for the signal case the $q^{2}$ dependence of the differential BR is a delta function with support at $q^{2}=m_{a}^{2}$, and the overall normalization $\mathcal{B}_{a}$ is factored out in Eq. (8). In the HTA case, one has to take into account that the invariant mass resolution of the experiment is finite. We do so by replacing $\delta(q^{2}-m_{a}^{2})$ with a normal distribution $g_{a}(q^{2})$ centered at the axion mass value with uncertainty taken to be $\sim(450~{\mathrm{M}eV})^{2}$, comparable to the beam-constrained mass resolution of $B$ mesons reconstructed using the hadronic tag [45]  [^6] In principle, the same smearing should also be applied to the di-neutrino channel, Eq. (12). In the axion case this treatment is essential, as it broadens the Dirac-delta shape of the $q^{2}$ distribution. For the di-neutrino mode, however, the impact is much milder because its $q^{2}$ spectrum is already smooth. Such an effect would become relevant only once a measurement of the differential $\mathcal{B}_{\nu\bar{\nu}}$ reaches comparable precision. In practice, the smeared and unsmeared $q^{2}$ distributions differ only marginally: a few events with small positive $q^{2}$ migrate to slightly negative reconstructed values, thus shifting between adjacent bins of the Belle II spectrum. Numerically, the neutrino likelihood at $\mathcal{B}_{a}=0$ remains nearly unchanged, and the bound on the axion is identical within the two-digit precision considered..
As regards the last component of Eq. (8), the numbers of expected background events for each bin, $n^{i}_{\mathrm{b}kg}$, can be read out from the plots of the collaboration. While the uncertainty on the Montecarlo generated background events are not explicitly provided, we infer that the uncertainty on the total normalisation is respectively $\sigma_{\text{ITA}}^{\text{bkg}}\approx 1\%$ and $\sigma_{\text{HTA}}^{\text{bkg}}\approx 10\%$. These estimates have been obtained from the $q^{2}$ distribution and its corresponding pull shown in Figs. 17 and 20 of Ref. [32]. One can in fact identify the pull with $|n^{i}_{\mathrm{d}at}-n^{i}_{\mathrm{h}st}|/(\sqrt{(\sigma^{i}_{\mathrm{d}at})^{2}+(\sigma^{i}_{\mathrm{h}st})^{2}})$, where the subscripts ‘dat’ and ‘hst’ denote respectively the data and the value of the stacked histograms, and $\sigma^{i}_{\mathrm{d}at}$ the uncertainty in the data point. Inverting the pull distribution in terms of $\sigma^{i}_{\mathrm{h}st}$ provides an estimate of the above uncertainty. Thus, we introduced two nuisance parameters, i.e. $\theta_{\text{ITA}}$ and $\theta_{\text{HTA}}$, both constrained by a normal Gaussian distribution in the likelihood as
${\mathcal{L}}_{2}=\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{\text{ITA}}/2}\times\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{\text{HTA}}/2}\,.$ (13)
As a non-trivial check of our procedure, we derive the profile likelihood ratio as a function of the signal strength $\mu_{\nu\bar{\nu}}$ in the case the axion is absent, i.e. ${\mathcal{B}}_{a}=0$, obtaining very consistent results as compared with the collaboration’s counterparts. Details are relegated to the text around Fig. 3 in the Supplemental Sec. A.
[…]
ITA HTA Combined
$\mu_{\nu\bar{\nu}}=1$ $n_{\text{dof}}=2$ $n_{\text{dof}}=2$ $n_{\text{dof}}=3$
$\mathcal{B}_{a}<1.1\times 10^{-6}$ $\mathcal{B}_{a}<5.1\times 10^{-6}$ $\mathcal{B}_{a}<1.4\times 10^{-6}$
$|(F_{V})_{sb}|>8.4\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>7.6\times 10^{8}$ GeV
$\mu_{\nu\bar{\nu}}\neq 1$ $n_{\text{dof}}=3$ $n_{\text{dof}}=3$ $n_{\text{dof}}=4$
$\mathcal{B}_{a}<1.0\times 10^{-6}$ $\mathcal{B}_{a}<6.2\times 10^{-6}$ $\mathcal{B}_{a}<1.2\times 10^{-6}$
$|(F_{V})_{sb}|>8.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.6\times 10^{8}$ GeV $|(F_{V})_{sb}|>8.0\times 10^{8}$ GeV
Table 1: Upper limits on the axion BR and lower limits on the coupling-rescaled Peccei-Quinn scale at 90% C.L. for the different cases considered. All limits assume $m_{a}=0$.
The different cases described and the resulting bounds are summarized in Table 1, which assumes $m_{a}=0$. Results for a general value of the axion mass are shown in Fig. 2. This figure quantitatively conveys two important pieces of information. First, as one would expect, the bound is weakened by an order of magnitude in the $m_{\nu\bar{\nu}}$ bins where data and the SM $B^{+}\to K^{+}\nu\bar{\nu}$ prediction are in tension. These discrepant bins can actually be used to set both an upper and a lower bound on $\mathcal{B}_{a}$ [47, 31, 42], as shown in Fig. 4 in the Supplemental Sec. A.
Second, and less obviously, the bound depends only marginally on the assumption of new physics in $\mathcal{B}_{\nu\bar{\nu}}$—namely, on whether $\mu_{\nu\bar{\nu}}$ is set to unity or left floating. This is an important conclusion, making the $B^{+}\to K^{+}\nu\bar{\nu}$ decay a double probe: of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, testable by $\mu_{\nu\bar{\nu}}\neq 1$, and of new light, elusive particles, produced as $B\to Ka$, with a BR sensitivity proportional to the total $\mathcal{B}_{\nu\bar{\nu}}$ uncertainty [48]. This second point is also conveyed by the similarity between the bounds in the first and second rows of Table 1. In the case where $\mathcal{B}_{\nu\bar{\nu}}$ is fixed to its SM value, we verify that varying it by as much as 10% has a negligible impact on these results. To corroborate this conclusion even further, we evaluate the bound on $\mathcal{B}_{a}$ assuming fixed values of $\mu_{\nu\bar{\nu}}$ but different from unity. As shown in Fig. 5 of the Supplemental Sec. A, the resulting bound has only a mild dependence on the specific value of the di-neutrino signal strength—even for huge beyond-SM contributions to the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude.
Figure 2: Upper limit on $\mathcal{B}_{a}$ (upper row) and lower limit on the coupling-rescaled Peccei-Quinn scale $(F_{V})_{sb}$ (lower row) at 90% CL as a function of the mass $m_{a}$, and with $\mu_{\nu\bar{\nu}}$ set to unity (blue line) or left floating (red).
The BR limit in Fig. 2 (upper row) can be translated into a bound on the coupling-rescaled Peccei-Quinn scale $(F_{V})_{sb}\equiv 2f_{a}/(k_{V})_{sb}$ [49], to obtain a bound on this scale. This bound is shown in Fig. 2 (lower row), as a function of $m_{a}$ for the two alternative $\mu_{\nu\bar{\nu}}$ possibilities. Limits for $m_{a}\simeq 0$ are collected in Table 1. As the table shows, we obtain $|(F_{V})_{sb}|\geq 0.9\times 10^{9}$ GeV for the QCD axion, which improves by one order of magnitude the pre-existing bound [49].
In conclusion, we reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of $B^{+}\to K^{+}a$, with $a$ an axion ($m_{a}\simeq 0$) or ALP ($m_{a}>0$). We derive analytically the mapping between the measured $q^{2}_{\text{rec}}$ and the true $q^{2}$, allowing the Inclusive Tag Analysis to constrain axion emission without relying on internal simulations. Using public efficiencies, we set the strongest existing limit on the Peccei–Quinn scale rescaled by the fundamental axion coupling relevant to our channel. The result is largely independent of the assumed $\mu_{\nu\bar{\nu}}$, making $B^{+}\to K^{+}\nu\bar{\nu}$ a dual probe of short-distance new physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude and of the production of an undetected axion. This approach can be generalized to $B^{+}\to K^{*+}\nu\bar{\nu}$—and to yet other channels—as soon as Belle II measurements become available.
Acknowledgments— We thank Joel C. Swallow for useful discussions. This work has received funding from the French ANR, under contracts ANR-19-CE31-0016 (‘GammaRare’) and ANR-23-CE31-0018 (‘InvISYble’), that we gratefully acknowledge.
A Supplementary plots
In this part we collect additional plots which are not necessary to understand the results in the main text, but further corroborate them, or can be used to perform comparisons.
Fig. 3 shows the log-likelihood profiles resulting from our analyses and their comparison with the Belle II counterparts. As the figure indicates, we obtain results very close to those of the collaboration, except in the region $\mu_{\nu\bar{\nu}}<1$ for the HTA case. This small deviation is understood: it originates from our treatment of the background uncertainty. When the background is fixed to its central value with no associated error, the minimum coincides with the Belle II result; including the background error, however, broadens the minimum and slightly lowers the best-fit point.
Fig. 4 shows $\mathcal{B}_{a}$ as a signal, displaying the $1\sigma$, $2\sigma$, and $3\sigma$ regions for $\mathcal{B}_{a}$ as a function of the mass, with $\mu_{\nu\bar{\nu}}$ either fixed to unity or left floating.
In Fig. 5, we evaluate the bound on $\mathcal{B}_{a}$ for fixed values of $\mu_{\nu\bar{\nu}}$ different from unity, illustrating the mild dependence of the resulting bound on the assumed value of the di-neutrino signal strength.
Figure 3: Twice the negative profile log-likelihood ratio as a function of $\mu_{\nu\bar{\nu}}$ for the ITA (red), HTA (blue) or combined (black) analyses. The solid line refers to our results while the dashed line are the results from Belle II in Fig. 16 of Ref. [32]. The grey dashed lines refer to $1,2,3\sigma$ $\chi^{2}$ values with 1 degree of freedom.
Figure 4: Regions $\mathcal{B}_{a}$ at 68% (green), 95% (yellow) and 99% (orange) CL and best fit value (solid green line) of $\mathcal{B}_{a}$ as a function of the mass with $\mu_{\nu\bar{\nu}}$ fixed to 1 (upper row) or left floating (lower row).
Figure 5: Upper limit at 90% CL on $\mathcal{B}_{a}$ for fixed values of $\mu_{\nu\bar{\nu}}$.